OpenStudy (anonymous):

Solve using the multiplication principle: -12x = 72

OpenStudy (anonymous):

x=-6

OpenStudy (tkhunny):

Please state the "Multiplication Principle".

OpenStudy (anonymous):

what r the options

OpenStudy (tkhunny):

Perhaps this? "Multiplying (or dividing) the same non-zero number to both sides of an equation does not change its solution set." Therefore, -12x = 72 Can be solved by dividing both sides by -12 -12x/(-12) = 72/(-12) x = -6

OpenStudy (anonymous):

thought my answer was wrong

OpenStudy (tkhunny):

@passive13 Your answer is wrong for two reasons: 1) It violates the standards of practice for OpenStudy, since you just dropped the answer with no explanation. 2) You simply solved it and failed to demonstrate how the "Multiplication Principle" was utilized in the solution - as required by the problem statement. Your answer is correct in that you have provided the solution.